A chemist titrates 230.0mL of a 0.4292M hydrocyanic acid HCN solution with 0.6266M KOH solution at 25°C. Calculate the pH at equivalence. The pKa of hydrocyanic acid is 9.21.
Round your answer to 2 decimal places.

Note for advanced students: you may assume the total volume of the solution equals the initial volume plus the volume of KOH solution added.
Answer #1

pH at equivalence = 11.31

Explanation

moles HCN = (concentration HCN) * (volume HCN)

moles HCN = (0.4292 M) * (230.0 mL)

moles HCN = 98.716 mmol

moles KOH required = moles HCN

moles KOH required = 98.716 mmol

volume KOH required = (moles KOH required) / (concentration KOH)

volume KOH required = (98.716 mmol) / (0.6266 M)

volume KOH required = 157.5 mL

Total volume at equivalence point = (volume HCN) + (volume KOH required)

Total volume at equivalence point = (230.0 mL) + (157.5 mL)

Total volume at equivalence point = 387.5 mL

At equivalence point, all HCN is neutralized to CN-

concentration CN- at equivalence = (moles HCN) / (Total volume at equivalence point)

concentration CN- at equivalence = (98.716 mmol) / (387.5 mL)

concentration CN- at equivalence = 0.2547 M

pKa = 9.21

pKb = 14 - pKa

pKb = 14 - 9.21

pKb = 4.79

Kb = 10-pKb

Kb = 10-4.79

Kb = 1.62 x 10-5

ICE table CN- (aq) H2O (l) HCN (aq) OH- (aq)
Initial conc. 0.2547 M - 0 0
Change -x - +x +x
Equilibrium conc. 0.2547 M - x - +x +x

Kb = [HCN]eq[OH-]eq / [CN-]eq

1.62 x 10-5 = [(x) * (x)] / (0.2547 M - x)

Solving for x, x = 2.02 x 10-3 M

[OH-] = x = 2.02 x 10-3 M

pOH = -log[OH-]

pOH = -log(2.02 x 10-3 M)

pOH = 2.69

pH = 14 - pOH

pH = 14 - 2.69

pH = 11.31